You are the owner of a restaurant and you just introduced three more items to your menu. Steak, shrimp and chicken platters sell for $55, $45, $25, respectively. Their variable costs are $15, $10, $5, respectively. You have a monthly fixed cost at $9,000. You have forecasted to sell 5,500 steaks, 7,000 shrimps and 10,000 Chicken platters annually.

1. Calculate your break-even point in average sales per day.

2. Calculate the number of steak, shrimp, and chicken platters that must be sold every day to break even.
Answer #1

Data:

selling price

variable cost

forecasted annual sales

Sales Proportion

monthly fixed cost

Steak

5

15

5500

0.244

-

Shrimp

45

10

7000

0.311

-

Chicken Platter

25

5

10000

0.444

-

22500

9000

Let the number of platters sold per day at break-even point is X.

No. of Steak Platters sold per day = 0.244 X

No. of Shrimp Platters sold per day = 0.311 X

No. of Chicken Platters sold per day = 0.444X

Total revenue = unit price * units sold

Total Revenue = 5*0.244 X + 45*0.311 X + 25*0.444 X = 26.32 X

Variable Cost = unit cost * units sold

Variable Cost = 15*0.244 X + 10*0.311 X + 5*0.444 X = 8.99 X

Monthly fixed cost = 9000

Per day = 9000 / 30 = 300

At breakeven point total revenue = total cost

26.32 X = 300 + 8.99 X

Or (26.32 - 8.99) * X = 300

Now X = 300 / 17.33 = 17.31 (approximately 17)

1. The breakeven point in average sales per day = 17

2. Replacing value of X:

No. of Steak Platters sold per day = 0.244 X = 0.244 * 17.31 = 4.22 or rounded to 4

No. of Shrimp Platters sold per day = 0.311 X = 0.311 * 17.31 = 5.38 or rounded to 5

No. of Chicken Platters sold per day = 0.444 X = 0.444 * 17.31 = 7.68 or rounded to 8

Approximatey 4, 5 and 8 platters of Steak, Shrimp and Chicken must be sold everyday to break-even.